Find three consecutive numbers where the product of the smaller two numbers is 37 less the square of the largest number.
Thank you!

Answer by stanbon(75887)

Find three consecutive numbers where the product of the smaller two numbers is 37 less the square of the largest number.
---
1st: x-1
2nd: x
3rd: x+1
---------------
Equation:
x(x-1) = (x+1)^2 - 37
---
x^2-x = x^2+2x+1 - 37
-x = 2x-36
3x = 36
x = 12
-------------
1st: x-1 = 11
2nd: x = 12
3rd: x+1 = 13
=====================
Cheers,
Stan H.